Two large flat plates are separated by a thin glycerin layer that is 5 mm thick and a temperature of 20° C. A force is applied to the top plate such shear stress = 26 Pa is generated. How fast will the top plate move relative to the bottom plate after a constant velocity is attained?
Answer #1

As we know the equation for torque

\\\tau = \mu \frac{V}{h} \\\\\text{Where,} \\\\\mu =\text{Viscosity of glycerin}=1.41 \ \frac{Ns}{m^2} \ \text{ At 20 degree and 100}\% \text{ Glycerin} \\\\V=\text{Velocity } \\\\\tau =\text{shear stress }=26 Pa \\\\h=\text{Thickness }=5mm=5*10^{-3} \ m

So,

\\v=\frac{\tau h}{\mu}=\frac{26*5*10^{-3}}{1.41} \\\\v=0.092198582 \ \ m/s

Hope it helpful to you for any doubt comment below i will solve your query
Rate my answer your rating is important to us

Table of the viscocity of glycerin at diff temp the unit of viscocity is \small \\\\\frac{mNs}{m^2}=\frac{10^{-3}Ns}{m^2}

Temperature (C)
Glycerine per cent weight 0 10 20 30 40 50 60 70 80 90 100
0 1.792 1.308 1.005 0.8007 0.656 0.5494 0.4688 0.4061 0.3565 0.3165 0.2838
10 2.44 1.74 1.31 1.03 0.826 0.68 0.575 0.5 0
20 3.44 2.41 1.76 1.35 1.07 0.879 0.731 0.635 0
30 5.14 3.49 2.5 1.87 1.46 1.16 0.956 0.816 0.69
40 8.25 5.37 3.72 2.72 2.07 1.62 1.3 1.09 0.918 0.763 0.668
50 14.6 9.01 6 4.21 3.1 2.37 1.86 1.53 1.25 1.05 0.91
60 29.9 17.4 10.8 7.19 5.08 3.76 2.85 2.29 1.84 1.52 1.28
65 45.7 25.3 15.2 9.85 6.8 4.89 3.66 2.91 2.28 1.86 1.55
67 55.5 29.9 17.7 11.3 7.73 5.5 4.09 3.23 2.5 2.03 1.68
70 76 38.8 22.5 14.1 9.4 6.61 4.86 3.78 2.9 2.34 1.93
75 132 65.2 35.5 21.2 13.6 9.25 6.61 5.01 3.8 3 2.43
80 255 116 60.1 33.9 20.8 13.6 9.42 6.94 5.13 4.03 3.18
85 540 223 109 58 33.5 21.2 14.2 10 7.28 5.52 4.24
90 1310 498 219 109 60 35.5 22.5 15.5 11 7.93 6
91 1590 592 259 127 68.1 39.8 25.1 17.1 11.9 8.62 6.4
92 1950 729 310 147 78.3 44.8 28 19 13.1 9.46 6.82
93 2400 860 367 172 89 51.5 31.6 21.2 14.4 10.3 7.54
94 2930 1040 437 202 105 58.4 35.4 23.6 15.8 11.2 8.19
95 3690 1270 523 237 121 67 39.9 26.4 17.5 12.4 9.08
96 4600 1580 624 281 142 77.8 45.4 29.7 19.6 13.6 10.1
97 5770 1950 765 340 166 88.9 51.9 33.6 21.9 15.1 10.9
98 7370 2460 939 409 196 104 59.8 38.5 24.8 17 12.2
99 9420 3090 1150 500 235 122 69.1 43.6 27.8 19 13.3
100 12070 3900 1410 612 284 142 81.3 50.6 31.9 21.3 14.8